OpenStudy (anonymous):

18.75 tons = ______________ lbs.

OpenStudy (anonymous):

18.75*2000=?

OpenStudy (mathstudent55):

How many pounds in a ton?

OpenStudy (anonymous):

2000 lbs in one ton

OpenStudy (anonymous):

1 ton equals 2000 lbs 18.75* 2000= 37,500lbs

OpenStudy (anonymous):

A ton is defined as different amounts in different parts of the world and in different contexts. Whichever value you're supposed to use for 1 ton, multiply it by 18.75.

OpenStudy (anonymous):

... its mostly standardized because he is asking lbs.

OpenStudy (anonymous):

37500

OpenStudy (anonymous):

yep

OpenStudy (mathstudent55):

Ok, since you know 1 ton = 2000 lb, then you know that to convert from tons to pounds you need to either multiply by 2000 or divide by 2000, right? How do you know which operation to do?

OpenStudy (anonymous):

from tons to pounds you multiply from pounds to tons you divide

OpenStudy (mathstudent55):

Since 1 ton = 2000 lb, that means a certain weight is a much higher number of pounds than tons. If you multiply 18.75 by 2000, you get a much higher number than 18.75. that's why you need to multiply.

OpenStudy (mathstudent55):

@Nogarenshi You are correct, of course, but the question is how do you get someone to figure it out if he forgets?

OpenStudy (anonymous):

for example they give you tons and you are asked to give pounds. Assuming you know the ratio 2000 lbs=1 ton. you can think that there are 2000 lbs in 1 ton. Ask yourself how many tons do the give you. Once you have that you know you have to multiply to get pounds \[tons*\frac{ lbs }{ tons }=lbs\]

OpenStudy (anonymous):

the tons cancel out leaving only lbs

OpenStudy (mathstudent55):

That, of course, is dimensional analysis, which is the best way of doing unit conversions.

OpenStudy (anonymous):

this is mainly something that you either have to memorize or figure out from the word play but otherwise dimensional analysis is the next best thing

OpenStudy (mathstudent55):

The reason dimensional anaylisis works is that since 1 ton = 2000 lb, dividing both sides by 1 ton, you get 1 = (2000 lb)/(1 ton), and dividing both sides by 2000 lb, you get (1 ton)/(2000 lb) = 1 Since the conversion factors equal 1, they do not change a number when you multiply by them. It onlt changes the units.

OpenStudy (anonymous):

exactly all it is is conversion of units nothing more and nothing less.

OpenStudy (anonymous):

so i think we pretty much covered this topic

OpenStudy (mathstudent55):

Yes, good job!

OpenStudy (anonymous):

hahaha you too

OpenStudy (anonymous):

even though this wasnt my question lol

OpenStudy (mathstudent55):

I know. Your good job was in explaining it.

OpenStudy (anonymous):

ohhh hahaha